A storage tank at STP contains 20.6 kg of nitrogen (N2). Volume of the tank is 16.4 m3 part b What is the pressure if an additional 33.4 kg of nitrogen is added without changing the temperature?
Answer #1

a) n = 20600 / 28 = 735.7 mol

V = n R T / P = (735.7 * 8.314 * 273.15) / (1.013 * 105)

=16.4 m3

b) n = 20600 + 33400 / 28 = 1928 mol

P = n R T / V = (1928 * 8.314 * 273.15) / (16.4)

= 2.67 * 105 Pa